# Exploring the Interplay between Musical Preferences and Personality through the Lens of Language

### v1

@@ 6.2 Statistical Significance of the Findings
We turn to evaluate the significance of differences between the groups’ scores. Our primary statistical test is ANOVA [14] – a statistical method that compares means across multiple groups to identify statistically significant differences. It analyzes both within-group and between-group variation to determine if observed differences are due to actual effects or random variation. ANOVA’s primary output, the F-statistic, indicates whether the means of the groups are significantly different. A $p$-value lower than 0.05 is commonly used as significance threshold. Applied on our per-community personality scores, the test results—all virtually zero, significant at pp<5.0e-46—strongly suggest that there are considerable differences in the means across the groups for each personality trait; the extremely low $p$-values confirm that these differences are highly significant.
@@ 6.2.1 Pairwise Significance and Effect Size
Statistical significance tests (group or pairwise) benefit from large samples under test, where even extremely small differences may show up significant. A common way to overcome this bias, is by reporting the additional measure of effect size. As such, Cohen’s $d$ effect size [9] quantifies the magnitude of the difference between two groups in terms of standard deviations, helping to understand the practical significance of the observed differences. A higher Cohen’s $d$ absolute value indicates a larger difference between the groups, where the common interpretation is as follows: absolute Cohen’s $d$ value between 0.2 and 0.5, denotes a small effect size, 0.5 to 0.8 – medium, and 0.8 or higher indicates high effect size.
@@
Figure 3: Cohen’s $d$ effect size and pairwise Mann-Whitney test. The value in each cell indicates the effect size, its color – whether the difference is significant at the $p$-val<0.05 level. Positive effect size (left group has a higher trait value) is marked with red, while negative effect size – with blue; both denote significant difference. Uncolored cells represent insignificant difference.
@@ 6.3 Predicting Genre from Personality
As an additional experiment, we approach the challenging task of predicting a user’s musical preference based solely on a five-dimensional personality vector derived from their authentic writing. Specifically, we trained another logistic regression classifier for five-class classification task: given a personality feature vector with five scores—OPN, AGR, CON, EXT, NEU—predict the individual’s musical preference out of the five genres we consider in this study. Notably, each user is associated in our study with a single genre (see Section 3). Applied on a random split of the 4,683 users into 80-20 train and test, the classifier achieved the accuracy of 0.424, considerably exceeding the five-class random baseline of 0.2: while insufficient for any practical purpose, this finding supports once again that personality traits "shine through" our language, and are associated with our musical preferences.
@@ 7 Conclusions
8 Ethical Considerations
We use publicly available data to study how individuals’ musical preferences are reflected in their spontaneous language through the lens of the Big Five personality traits. The use of publicly available data from social media platforms, such as Reddit, may raise normative and ethical concerns. Here we address two main concerns:
Anonymity: Data used for this research can only be associated with participants’ user IDs, which, in turn, cannot be linked to any identifiable information, or used to infer any personal or demographic trait. Consent: Jagfeld et al. [18] debated the need to obtain informed consent for using social media data mainly because it is not straightforward to determine if posts pertain to a public or private context. Ethical guidelines for social media research [5] and practice in comparable research projects [2], as well as Reddit’s terms of use, [^14] https://www.redditinc.com/policies/user-agreement-september-12-2021 regard it as acceptable to waive explicit consent if users’ anonymity is protected.
@@ 8 Ethical Considerations
9 Limitations
Despite our efforts to carefully mitigate potential confounds, this study is not without limitations. The use of Reddit as a data source introduces certain biases, as its demographic skews toward younger males, which may affect the generalizability of our findings. Additionally, assuming that the most active participants in genre-specific subreddits are representative of the broader fan base is a strong, though not unreasonable, assumption, as not all subscribers may strongly identify with the genre. Moreover, (unknown) factors such as age and gender were not explicitly controlled for, potentially influencing the results.
Nevertheless, the large scale of our dataset helps enhance generalizability and mitigate demographic biases. Furthermore, prior research has validated the approach of using high activity in topic-related subreddits as a proxy for user affiliation with that topic.
@@ A.1 Personality Traits (Short) Definitions
Openness (also referred to as Openness to experience) emphasizes imagination and insight the most out of all five personality traits. People who are high in Openness tend to have a broad range of interests. They are curious about the world and other people and are eager to learn new things and enjoy new experiences. People who are high in this personality trait also tend to be more adventurous and creative. Conversely, people low in this personality trait are often much more traditional and may struggle with abstract thinking. Intellect, imagination, and Openness describe your imagination and how creative you are. It refers to your sense of curiosity about the world and your willingness to try new things.

### v2

@@ 6.3 Statistical Significance of the Findings
We turn to evaluate the significance of differences between the groups’ scores. Our primary statistical test is ANOVA [32] – a statistical method that compares means across multiple groups to identify statistically significant differences. It analyzes both within-group and between-group variation to determine if observed differences are due to actual effects or random variation. Particularly useful for experiments involving more than two groups, ANOVA extends the capabilities of the two-group t-test. ANOVA’s primary output, the F-statistic, indicates whether the means of the groups are significantly different. A $p$-value lower than 0.05 is typically used as significance threshold.
When applied on our per-community personality scores across each dimension, the test results — all virtually zero, significant at pp<5.0e-46 — strongly suggest that there are considerable differences in the means across the groups for each personality trait; the extremely low $p$-values confirm that these differences are significant.
@@ 6.3.1 Pairwise Significance and Effect Size
Statistical significance tests benefit from large samples under test, where even extremely small differences may show up significant. A common way to overcome this bias, is by reporting the additional measure of effect-size. As such, Cohen’s $d$ effect size [34] quantifies the magnitude of the difference between two groups in terms of standard deviations, helping to understand the practical significance of the observed differences. A higher Cohen’s $d$ absolute value indicates a larger difference between the groups, where the common interpretation is as follows: absolute Cohen’s $d$ value between 0.2 and 0.5, denotes a small effect size, 0.5–0.8 – medium, and 0.8 or higher indicates high effect size.
@@
Figure 3: Cohen’s $d$ effect size and pairwise Mann-Whitney test. The value in each cell indicates the effect size, its color – whether the difference is significant at the $p$-val<0.05 level. Positive effect size (left group has a higher trait value than right) is marked with red, while negative effect size – with blue; both denote significant difference. Uncolored cells represent insignificant difference.
@@ 6.4 Predicting Genre from Personality
As an additional experiment, we approach the challenging task of predicting a user’s musical preference based solely on a five-dimensional personality vector derived from their authentic writing. Specifically, we trained another logistic regression classifier for five-class classification task: given a personality feature vector with five scores — OPN, AGR, CON, EXT, NEU — predict the individual’s musical preference out of the five genres we consider in this study. Notably, each user is associated in our study with a single genre (see Section 3). Applied on a random split of the 4,683 users into 80-20 train and test, the classifier achieved the accuracy of 0.424, considerably exceeding the five-class random baseline of 0.2: while insufficient for any practical purpose, this finding supports once again that personality traits "shine through" our language, and here, as well, are associated with our musical preferences.
@@ 7 Conclusions
Ethical Considerations
We use publicly available data to study how individuals’ musical preferences are reflected in their spontaneous language through the lens of the Big Five personality traits. The use of publicly available data from social media platforms, such as Reddit, may raise normative and ethical concerns. These concerns are extensively studied by the research community as reported in e.g., Proferes et al. [35].
Here we address two main concerns: Anonymity: Data used for this research can only be associated with participants’ user IDs, which, in turn, cannot be linked to any identifiable information, or used to infer any demographic trait. Consent: Jagfeld et al. [36] debated the need to obtain informed consent for using social media data mainly because it is not straightforward to determine if posts pertain to a public or private context. Ethical guidelines for social media research [37] and practice in comparable research projects [38], as well as Reddit’s terms of use, [^14] https://www.redditinc.com/policies/user-agreement-september-12-2021 regard it as acceptable to waive explicit consent if users’ anonymity is protected.
@@
Limitations
Despite our efforts to carefully mitigate potential confounds, this study is not without limitations. The use of Reddit as a data source introduces certain biases, as its demographic skews toward younger males, which may affect the generalizability of our findings. Additionally, assuming that the most active participants in genre-specific subreddits are representative of the broader fan base is a strong, though not unreasonable, assumption, as not all subscribers may strongly identify with the genre. Moreover, (unknown) factors such as age and gender were not explicitly controlled for, potentially influencing the results. Nevertheless, we believe that the large scale of our dataset helps enhance generalizability and mitigate demographic biases. Furthermore, prior research has validated the approach of using high activity in topic-related subreddits as a proxy for user affiliation with that topic.
Acknowledgements.
We are grateful to our four anonymous reviewers for their useful comments and constructive feedback.
Declaration on Generative AI
The authors have not employed any Generative AI tools while writing the paper.
@@ A.1 Personality Traits (Short) Definitions
Openness (also referred to as Openness to experience) emphasizes imagination and insight the most out of all five personality traits. People who are high in Openness tend to have a broad range of interests. They are curious about the world and other people and are eager to learn new things and enjoy new experiences. People who are high in this personality trait also tend to be more adventurous and creative. Conversely, people low in this personality trait are often much more traditional and may struggle with abstract thinking. Intellect, imagination, and Openness describe your imagination and how creative you are. It refers to your sense of curiosity about the world and your willingness to try new things, and to be exposed to new experiences.